# Towards Robust Real-World Multivariate Time Series Forecasting: A Unified Framework for Dependency, Asynchrony, and Missingness

### v1

@@ 4 ChannelTokenFormer
To jointly model temporal patterns within each channel and cross-channel dependencies, we design a unified masked self-attention mechanism that integrates both local and global representations into a single attention operation (see Figure 2). Each channel $i\in\{1,\dots,N\}$ contributes two types of tokens: (i) local tokens $\mathbf{L}^{(i)}\in\mathbb{R}^{L^{(i)}\times d}$ representing patch-level embeddings, and (ii) global channel tokens $\mathbf{C}^{(i)}\in\mathbb{R}^{C\times d}$ summarizing high-level contextual information, where $L^{(i)}$ and $C$ denote the number of local and channel tokens, respectively. The full token sequence is constructed as $\mathbf{X}=[\mathbf{L}^{(1)};\mathbf{C}^{(1)};\dots;\mathbf{L}^{(N)};\mathbf{C}^{(N)}]\in\mathbb{R}^{T\times d}$, with total token count $T=\sum_{i=1}^{N}(L^{(i)}+C)$. This unified sequence is passed through a masked multi-head self-attention layer with residual connection: $\mathbf{X}_{\text{out}}=\mathbf{X}+\text{Attention}(Q,K,V)$, where $Q=\mathbf{X}\mathbf{W}_{Q}$, $K=\mathbf{X}\mathbf{W}_{K}$, and $V=\mathbf{X}\mathbf{W}_{V}$ are linear projections of $\mathbf{X}$, and the attention is computed as $\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+\mathbf{M}\right)V$. Here, $\mathbf{M}\in\mathbb{R}^{T\times T}$ encodes structural constraints on how tokens can attend to one another, reflecting their types, channels, and masking. Our tailored masking scheme defines the construction of $\mathbf{M}$ as follows (See Figure 3): (1) Local tokens attend only to other local tokens within the same channel, enabling intra-temporal modeling. (2) Channel tokens attend to their own local tokens and to other channels’ tokens, but are not accessible to local tokens due to their read-only role in the attention mechanism. (3) Channel tokens do not attend to themselves; for each query $\mathbf{C}^{(i)}$, the key $\mathbf{C}^{(i)}$ is excluded to avoid self-reinforcement and to encourage informative cross-channel interaction. Our unified masking strategy, inspired by read-only prompt attention [21] and attention control techniques [19], is implemented without modifying the standard Transformer architecture. Stacked masked attention blocks with residual connections and layer normalization enable deeper modeling of temporal and cross-channel dependencies.
While the channel-wise design of CTF can lead to parameter overhead, unifying local and cross-channel attention into a single layer helps offset this cost and improves parameter efficiency. This design provides an inductive bias that simultaneously supports temporal abstraction, captures channel dependencies, and enhances robustness to missing intervals and asynchronous sampling—key challenges in real-world forecasting.
@@ 5 Experiments
To assess the effectiveness of ChannelTokenFormer under the newly proposed problem settings, we conduct extensive experiments on three multivariate time series benchmark datasets that we modify to reflect practical conditions, and on one real-world dataset from the LNG cargo handling system of an LNG carrier.
@@
We evaluate forecasting performance under channel-wise asynchronous sampling using four multivariate datasets. ETT1-practical (ETT1) modifies ETTm1 [55] by resampling channels to domain-specific temporal resolutions (e.g., 1-hour for load, 15-minute for temperature). Weather-practical (Weather), adapted from the Weather benchmark [43], assigns heterogeneous sampling periods to channels based on distinct physical characteristics. Monash-SolarWindPower-practical (SolarWind) combines solar and wind power series from the Monash Forecasting Archive [10], with variables resampled to 20-minute and 5-minute intervals, respectively. LNG Cargo Handling System (CHS) is a real-world industrial dataset collected from an LNG carrier, consisting of sensor channels related to cargo operations, ship navigation, and surrounding weather conditions. To assess robustness under block-wise test-time missing intervals, we additionally conduct experiments on SolarWind by varying the missing ratio. Detailed dataset specifications are provided in Appendix A.1.
@@
To provide a comprehensive evaluation, we compare our CTF against representative state-of-the-art models from diverse architectural paradigms. We include Transformer-based models such as TimeXer [42], iTransformer [29], and PatchTST [34], along with the CNN-based TimesNet [44], the GNN-based CrossGNN [15], and MLP-based models TimeMixer [39] and DLinear [50].
@@
Most baseline models assume regularly sampled and fully observed inputs. To meet this assumption, we use linear interpolation or forward filling to recover values unobserved due to differences in channel-specific sampling periods. However, to simulate test-time block-wise missingness, we randomly select contiguous regions in the input to be masked and apply zero-filling instead of other interpolation methods. This decouples the impact of missingness from asynchronous sampling and creates a conservative evaluation setting that amplifies differences between models with and without explicit missing-aware mechanisms. For each dataset, input and prediction lengths are selected to reflect the channel-wise temporal resolution. To accommodate heterogeneous sampling periods across channels, we define these lengths based on a fixed time duration rather than a fixed number of steps. The actual number of input and output points per channel is determined by its sampling period, with shorter periods yielding more points over the same window. Detailed settings are provided in Appendix A.2.
@@ 5.1 Main Results
Scenario 1: Channel-wise Asynchronous Forecasting
We evaluate forecasting performance without test-time missing intervals using our newly constructed datasets that reflect realistic sampling heterogeneity across channels. Unlike baselines that rely on interpolation, our approach preserves the original sampling structure. As summarized by average CMSE and CMAE for all prediction lengths in Table 1, our method consistently outperforms all baselines across datasets in the practical setting. These results highlight the advantage of avoiding interpolation in asynchronous environments. Detailed results for each prediction length and dataset are provided in Appendix E.
Table 1: (Scenario 1) Forecasting performance on the channel-wise asynchronous forecasting task.
Strategy
@@
0.286
0.141
0.334
0.141
0.316
0.132
0.338
0.236
Scenario 2: Channel-wise Asynchronous Forecasting with Block-wise Test-time Missing
We further evaluate all models under a more challenging setting, where block-wise missing intervals are applied to test-time inputs, on top of channel-wise asynchronous sampling. In this scenario, all models receive inputs containing contiguous missing segments filled with zeros. Although zero-filling is applied for input compatibility, our model uniquely utilizes masking to prevent attention to regions treated as missing. As shown in Table 2, our model consistently outperforms all baselines in average CMSE and CMAE for all prediction lengths across different missing ratios, maintaining top performance even as the severity of missingness increases. Notably, at a missing ratio of 0.25, the performance degradation of our model remains as low as 1.23%, indicating its strong resilience to partially missing inputs. This robustness stems from the use of random patch masking during training, which improves the model’s ability to generalize to incomplete patterns. Unlike other methods, our model avoids distortion from artificially zero-filled values and preserves signal fidelity under block-wise missing conditions. Detailed results across datasets are provided in Appendix E.
Table 2: (Scenario 2) Forecasting performance on the SolarWind dataset under block-wise test-time missingness with varying missing ratios $m$.
Strategy
@@
$m=0.0$
0.405
0.455
0.420
@@ 5.2 Ablation Study
Table 3: Ablation study under channel-wise asynchronous sampling with block-wise test-time missing (missing ratio = 0.25). Results are averaged across all prediction lengths on the SolarWind dataset.
@@
w/o Channel Dependence (CI only) 0.435 0.484
w/o Dynamic patching 0.430 0.492
w/o Patch masking 0.486 0.525
To assess the effectiveness of our design in addressing the three key challenges—Dependency, Asynchrony, and Missingness—we conduct a structured ablation study in which each component of our architecture is selectively removed or replaced. For Dependency, we disable cross-channel attention by restricting interactions between channel tokens from different channels. In the case of Asynchrony, we remove the dynamic patching mechanism, forcing the model to process input sequences with varying lengths using fixed patch boundaries that ignore dominant frequencies and channel-specific sampling periods. For Missingness, we eliminate patch masking during both training and testing, thereby exposing the model to inputs with artificially zero-filled values in place of structured missingness handling. Robust forecasting in practical settings requires all three components—channel dependence, dynamic patching, and patch masking—as the removal of any one leads to degraded performance. Additional ablations are provided in Appendix C.1.
@@ 6 Analysis
6.1 Effect of Number of Channel Tokens
To examine the impact of channel token granularity, we vary the number of channel tokens across all practical datasets. As shown in Table 5, the optimal channel token count differs by dataset, reflecting variations in size, temporal complexity, and signal diversity. Two tokens per channel yield the best performance on ETT1 and SolarWind, balancing expressiveness and generalization. Weather benefits from more tokens due to its multi-scale, high-variability signals, whereas CHS performs best with a single token, due to sparse, low-frequency control signals where over-tokenization leads to overfitting. These results suggest tuning the number of channel tokens based on dataset characteristics—favoring fewer tokens for small or redundant datasets, and more for rich, dynamic ones.
@@ 6.1 Effect of Number of Channel Tokens
# Channel Tokens ETT1 SolarWind Weather CHS
@@ 6.2 Robustness to Input Length Variability
Unlike conventional forecasting models restricted to fixed-length inputs, our method is trained with random patch masking to improve robustness to variable-length sequences. As shown in Table 5, the model maintains stable performance across all test input lengths on the SolarWind dataset without requiring retraining. Since each channel has a different sampling period, we define the input length based on the channel with the smallest sampling interval, which corresponds to 576 time steps over 2 days. Even when this length is reduced to 288 (equivalent to 1 day), the CMSE increases by only 4.9%, demonstrating strong generalization to input length variation, which is an essential property for real-world forecasting under inconsistent input conditions.
@@ 6.3 Frequency Bias Analysis in Predicted Time Series
We evaluate the impact of interpolation on the frequency characteristics of predicted time series under practical conditions with channel-wise asynchronous sampling periods. To this end, we compare the spectral fidelity of forecasts from our interpolation-free model, ChannelTokenFormer (CTF), with that of TimeXer, a representative interpolation-based baseline. The analysis focuses on dominant frequency difference and amplitude spectrum RMSE computed across low, mid, and high frequency bands, with Table 6 reporting overall means averaged across all channels in the ETT1 dataset. Zero-centering is applied prior to spectral analysis to remove ultra-low-frequency (DC) components and enhance periodicity detection. In the ETT1 dataset, dominant periodicities are concentrated in the low- to mid-frequency range (e.g., 8, 12, and 24 hours), where interpolation-induced phase delay has limited effect. As a result, even TimeXer, which relies on interpolation, is able to localize the dominant spectral peak with an accuracy comparable to that of CTF. However, the amplitude spectrum RMSE reveals consistent discrepancies across all frequency bands, indicating that CTF more effectively preserves the magnitude of spectral components. As illustrated in Figure 4, TimeXer undergoes greater amplitude attenuation compared to CTF. In contrast, CTF better maintains the original spectral energy. These results indicate that CTF mitigates spectral distortion introduced by interpolation and preserves frequency structure under asynchronous sampling, which translates to improved trend and periodicity retention in the time domain.
Table 6: Frequency-domain comparison between CTF (interpolation-free) and TimeXer (interpolation-based) on the ETT1 dataset for a prediction length of 192 (2-day window). Lower values indicate better spectral fidelity. RMSEs are computed over three frequency bands under 1-hour sampling ($f_{s}=1$): Low (0.00–0.10), Mid (0.10–0.25), and High (0.25–0.50) cycles/hour, corresponding to periodicities longer than 10 hours, between 4 and 10 hours, and shorter than 4 hours, respectively.
@@ 7 Conclusion
We introduce a practical forecasting setting that reflects real-world conditions, including channel-wise asynchronous sampling and block-wise missing intervals. Instead of relying on interpolation, our approach handles both missing segments and sampling gaps through masking and frequency-based dynamic patching. Built on this setting, our unified attention model leverages channel tokens with read-only masking and cross-channel attention, consistently outperforming prior methods across real-world scenarios. However, as the number of channels increases, full attention across all channels, based purely on masking to regulate information flow, may scale less efficiently. To support high-dimensional forecasting, future work will focus on developing more scalable versions of our model.
Appendix

### v2

@@ A Unified Mask-Guided Attention with Channel Tokens for Real-World Challenges
To jointly address three practical challenges in multivariate time series, we design a unified, mask-guided self-attention mechanism that integrates local and global representations within a single attention operation. As illustrated in Figure 2, this design unifies both token types in one attention step. Each channel $i\in\{1,\dots,N\}$ contributes two types of tokens: (i) local tokens $\mathbf{L}^{(i)}\in\mathbb{R}^{L^{(i)}\times d}$ representing patch-level embeddings, and (ii) global channel tokens $\mathbf{C}^{(i)}\in\mathbb{R}^{C\times d}$ summarizing high-level contextual information, where $L^{(i)}$ and $C$ denote the number of local and channel tokens, respectively. The full token sequence is constructed as $\mathbf{X}=[\mathbf{L}^{(1)};\mathbf{C}^{(1)};\dots;\mathbf{L}^{(N)};\mathbf{C}^{(N)}]\in\mathbb{R}^{T\times d}$, with total token count $T=\sum_{i=1}^{N}(L^{(i)}+C)$. This unified sequence is passed through a masked multi-head self-attention layer with residual connection: $\mathbf{X}_{\text{out}}=\mathbf{X}+\text{Attention}(Q,K,V)$, where $Q=\mathbf{X}\mathbf{W}_{Q}$, $K=\mathbf{X}\mathbf{W}_{K}$, and $V=\mathbf{X}\mathbf{W}_{V}$ are linear projections of $\mathbf{X}$, and the attention is computed as $\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+\mathbf{M}\right)V$. Here, $\mathbf{M}\in\mathbb{R}^{T\times T}$ encodes structural constraints on how tokens can attend to one another, reflecting their types, channels, and masking. Our tailored masking scheme constructs $\mathbf{M}$ as illustrated in Figure 3: (1) Local tokens attend only to other local tokens within the same channel, enabling intra-temporal modeling. (2) Channel tokens attend to their own local tokens and to other channels’ tokens, but are not accessible to local tokens due to their read-only role in the attention mechanism. (3) Channel tokens do not attend to themselves; for each query $\mathbf{C}^{(i)}$, the key $\mathbf{C}^{(i)}$ is excluded to avoid self-reinforcement and to encourage informative cross-channel interaction. Our unified masking strategy, inspired by read-only prompt attention (Lee et al., 2023) and attention control techniques (Kim et al., 2025), is implemented without modifying the standard Transformer architecture. Stacked masked attention blocks with residual connections and layer normalization enable deeper modeling of temporal and cross-channel dependencies. Although channel-wise tokenization introduces additional parameters, the masking strategy ensures computation remains efficient by restricting attention to necessary token interactions, while scaling reliably to realistic channel counts and sequence lengths, as shown in Appendix C.3.
@@ 5 Experiments
To assess the effectiveness of ChannelTokenFormer under our proposed problem settings, we conduct extensive experiments on four widely used multivariate time series benchmarks adapted to practical conditions, and on two real-world datasets from air monitoring and the LNG cargo handling system. Additional results under conventional settings are provided in Appendix D.
@@
We evaluate forecasting performance under channel-wise asynchronous sampling using four multivariate datasets. ETT1-practical and ETT2-practical (ETT1 & ETT2) modify ETTm1 and ETTm2 (Zhou et al., 2021) by resampling channels to domain-specific temporal resolutions (e.g., 1-hour for load, 15-minute for temperature). Weather-practical (Weather), adapted from the Weather benchmark (Wu et al., 2021), assigns heterogeneous sampling periods to channels based on distinct physical characteristics. Monash-SolarWindPower-practical (SolarWind) combines solar and wind power series from the Monash Forecasting Archive (Godahewa et al., 2021), with variables resampled to 20-minute and 5-minute intervals, respectively. EPA-Air (EPA), adapted from the U.S. Environmental Protection Agency’s air quality monitoring data (Chang et al., 2025), collects measurements from several regions with heterogeneous sampling periods across channels (e.g., 1-hour for temperature, 8-hour for PM2.5, among others). We choose four regions, Maricopa, Richmond, LA and Hillsborough. LNG Cargo Handling System (CHS) is a real-world industrial dataset collected from an LNG carrier, consisting of sensor channels related to cargo operations, ship navigation, and surrounding weather conditions. To assess robustness under block-wise test-time missing intervals, we additionally conduct experiments on SolarWind by varying the missing ratio. Detailed dataset specifications are provided in Appendix A.1.
@@
We compare CTF against representative state-of-the-art models across architectures and methods for irregularity and missingness. Transformer baselines include TimeXer (Wang et al., 2024e), iTransformer (Liu et al., 2024c), and PatchTST (Nie et al., 2023). We also cover CNNs, GNNs, and MLPs, including TimesNet (Wu et al., 2023), CrossGNN (Huang et al., 2023), TimeMixer++ (Wang et al., 2024a), and DLinear (Zeng et al., 2023). We further include irregular-sampling methods such as ContiFormer (Chen et al., 2023), t-PatchGNN (Zhang et al., 2024), and Hi-Patch (Luo et al., 2025) as well as a missingness-robust approach, BiTGraph (Chen et al., 2024b).
@@
Baseline models for regular time series assume uniformly sampled and fully observed inputs. To satisfy this requirement, we employ linear interpolation to recover values unobserved at timestamps not covered by channel-specific sampling periods. To isolate architectural effects from interpolation, we further implement interpolation-free variants, including a version of TimeXer (Wang et al., 2024e) modified to handle temporally non-aligned cross-channel inputs. Implementation details are provided in Appendix C.4. In contrast, irregular or missing-aware baselines do not require interpolation, as they operate solely on observed values. To simulate block-wise missingness at test time, we randomly mask contiguous input regions. For fairness, regular baselines without explicit missing-value handling are also provided with linearly interpolated values for these masked regions, consistent with their training procedure. For each dataset, input and prediction lengths are selected to reflect the channel-wise temporal resolution. Despite heterogeneous sampling periods across channels, it is practical to define input and prediction windows by a fixed time duration instead of requiring the same counts of sample points across channels. The actual number of input and output points per channel is determined by its sampling period, with shorter periods yielding more points over the same window.
@@ 5.1 Main Results
Case 1: Channel-wise Asynchronous Forecasting
We evaluate forecasting performance using datasets that reflect realistic sampling heterogeneity across channels. Unlike regular baselines that rely on interpolation, our approach preserves the original sampling structure and operates directly on observed values. Averaging CMSE and CMAE over all prediction lengths (Table 1), our method achieves the best performance on all datasets in the practical setting, highlighting the benefit of avoiding imputation when channels are temporally misaligned. Our method also handles asynchrony more effectively than irregular baselines. Detailed results for each prediction length and dataset are provided in Appendix E.
Table 1: Forecasting performance on the channel-wise asynchrony (Case 1).
Approach
@@
0.351
0.248
0.301
0.129
0.294
0.125
Case 2: Channel-wise Asynchronous Forecasting with Test-time Missing Blocks
We further evaluate all models under a more challenging setting, where test-time inputs include block-wise missing intervals, explicitly indicated to the model, on top of channel-wise asynchronous sampling. This configuration encompasses all three core challenges of interest. In this case, all models receive inputs containing contiguous missing segments filled with zeros. But for a fair comparison, linear interpolation is applied to regular baselines to avoid underestimating their capability. Unlike other approaches, our model can internally adjust input length by design, enabling the use of zero-patch masking to effectively handle missing regions. As shown in Table 2, our model consistently outperforms all baselines in average CMSE and CMAE for all prediction lengths across different missing ratios, maintaining high performance even as the severity of missingness increases. This robustness stems from the use of random patch masking during training, which improves the model’s ability to generalize to incomplete patterns. Unlike other methods, our model avoids distortion from artificially zero-filled values and preserves signal fidelity under block-wise missing conditions. Detailed results across datasets are provided in Appendix E. To further demonstrate the overall robustness to missing handling, we conducted additional experiments on the ETT1 and EPA datasets with smaller interval missing settings. The corresponding details are provided in Appendix A.2.
Table 2: Channel-wise Asynchronous Forecasting performance on the SolarWind dataset under block-wise test-time missingness with varying missing ratios $m$ (Case 2).
Approach
@@
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
$m=0.125$
0.409
0.463
0.427
0.474
0.475
0.491
0.472
0.488
@@ 5.2 Ablation Study
Table 3: Ablation study under channel-wise asynchronous sampling with block-wise test-time missing (missing ratio = 0.375). Results are averaged across all prediction lengths on the SolarWind dataset.
@@
w/o Channel Dependence (CI only) 0.474 0.521
w/o Dynamic patching 0.494 0.536
w/o Patch masking 0.458 0.508
To assess the effectiveness of our design in addressing the three key challenges, we conduct an ablation study where each component is selectively removed or replaced. For Dependency, we disable cross-channel attention. For Asynchrony, we remove the dynamic patching mechanism, forcing fixed patch boundaries that ignore dominant frequencies and channel-specific sampling periods. For Missingness, we eliminate patch masking during training and test time, exposing the model to zero-filled inputs instead. These experiments show that robust forecasting under practical conditions requires all three components, since removing any one degrades performance. Additional ablations are provided in Appendix C.1.
@@ 6 Analysis
6.1 Understanding the Optimal Number of Channel Tokens
The number of channel tokens was varied to analyze its impact on asynchronous forecasting. Results in Table 5 show that the optimal choice differs by dataset: two channel tokens work best for ETT1 and SolarWind, three for EPA, and one for Weather and CHS. These outcomes reflect how channels interact under asynchronous sampling. When channels move in a similar and highly correlated way, a single channel token captures the shared pattern and more tokens add redundancy. ETT1 and SolarWind, which exhibit two dominant groups of channels, benefit from two tokens. EPA shows stronger heterogeneity, and three tokens better capture its diverse structures. Overall, the optimal number of channel tokens depends on the degree of channel similarity and diversity in each dataset.
@@ 6.1 Understanding the Optimal Number of Channel Tokens
Dataset
# Channel Tokens ETT1 ETT2 SolarWind Weather EPA CHS
@@ 6.2 Robustness to Input Length Variability
Random patch masking is primarily designed to make the model robust to incomplete inputs caused by block-wise missing intervals at test time. An additional benefit of this design is robustness to input length variability, which remains a limitation for conventional forecasting models. Most existing approaches train and evaluate separate models for each target input length, and even models that technically accept variable lengths often degrade in performance when the sequence length changes. As shown in Table 5, CTF maintains stable performance across all test input lengths on the SolarWind dataset without requiring retraining. Since each channel has a different sampling period, we define the input length based on the channel with the smallest sampling interval, which corresponds to 576 time steps over 2 days. Even when this length is reduced to 288 (equivalent to 1 day), the CMSE increases by only 4.9%, demonstrating strong generalization to input length variation, which is another essential property for real-world forecasting under inconsistent input conditions.
@@ 6.3 Frequency Bias Analysis in Predicted Time Series
We analyze how interpolation affects the frequency characteristics of predicted time series under practical conditions with channel-wise asynchronous sampling. To this end, we focus on TimeXer (Wang et al., 2024e), which is structurally similar to CTF and achieves nearly identical performance to ours in the conventional setting on the ETTm1 benchmark. This makes TimeXer a strong reference point for fair comparison. In the practical setting, however, TimeXer can only process asynchronous inputs by performing interpolation, since its original design cannot directly accommodate non-aligned sequences without architectural modifications. As a result, any performance gap observed here reflects interpolation-induced distortion rather than inherent modeling differences. We analyze dominant frequency difference and amplitude spectrum RMSE across frequency bands in the ETT1 dataset (Table 6). In ETT1 (Table 6), both models localize dominant cycles similarly, since the main periodicities (e.g., 8, 12, and 24 hours) lie in the low- to mid-frequency range where interpolation-induced phase delay is minimal. However, amplitude spectrum RMSE shows consistent gaps: as Figure 4 illustrates, both models exhibit some attenuation, but the effect is noticeably stronger for TimeXer, while CTF better preserves spectral energy. Thus, although TimeXer matches CTF in conventional settings, its reliance on interpolation in the practical setting introduces frequency bias that propagates into predictions. By avoiding interpolation, CTF mitigates this distortion and better retains trend and periodicity under asynchronous sampling.
Table 6: Frequency-domain comparison between CTF (interpolation-free) and TimeXer (interpolation-based) on the ETT1 for a prediction length of 192 (2-day window). Lower values indicate better spectral fidelity. RMSEs are computed over three frequency bands under 1-hour sampling ($f_{s}=1$): Low (0.00–0.10), Mid (0.10–0.25), and High (0.25–0.50) cycles/hour, corresponding to periodicities longer than 10 hours, between 4 and 10 hours, and shorter than 4 hours, respectively.
@@ 7 Conclusion
A practical forecasting setting is introduced where three core real-world challenges occur simultaneously: channel-wise asynchronous sampling, block-wise missing intervals, and complex cross-channel dependencies. Instead of interpolation, our approach handles both missing segments and sampling gaps through masking and frequency-based dynamic patching. The unified mask-guided attention leverages channel tokens with read-only masking and cross-channel attention, consistently outperforming prior methods across real-world conditions. Scalability analysis further shows that the masking-based design remains efficient for realistic channel counts and input lengths, although extreme scales may still pose challenges. Beyond scaling, future work will also consider applying ChannelTokenFormer to broader domains, integrating multimodal signals, and improving interpretability of channel interactions, thereby expanding its utility in real-world forecasting.